You have landed on an unknown planet, Newtonia, and want to know what objects will weigh there. You find that when a certain tool is pushed on a frictionless horizontal surface by a 12.5 N force, it moves 16.4 m in the first 2.40 s , starting from rest. You next observe that if you release this tool from rest at 10.1 m above the ground, it takes 2.88 s to reach the ground.
Answer #1

here,

for the tool , let the mass is m

when it is pushed horizontally ,

using second equation of motion

d= 0.5 at^2

16.4 = 0.5 a * 2.4^2

a = 5.69 m/s^2

Using second law of motion

F =m*a

12.5 = 5.69 * m

m = 2.195 Kg

Now, when it is thrown vertically

as h = .5 * g * t^2

10.1 = 0.5 * g * 2.88^2

g = 2.44 m/s^2

Now , weight of object = m*g

weight of object = 2.44 * 2.195

weight of object = 5.35 N

the weight of object is 5.35 N